If the statement is true for all integer values of p and q,
explain why. If it isn’t true, give a counterexample.
a. /p/is positive. ( with straight lines though not /)
b. -q is negative.
c. If p is the opposite of q, p+q=0
Thank you
Answer by richard1234(7193) 
 a: |p| is positive is not always true. Note that p could equal 0, and |0| = 0, which is not positive.
b: -q is negative is also not always true, since q could be negative. For example, if q = -1, then -q = -(-1) = 1.
c: This must always be true. If p = -q, then p + q = (-q) + q = 0 (for all values p,q with p = -q)
| |
|